The table to the right shows the results of a survey in which

26002600

adults from Country​ A,

11081108

adults from Country​ B, and

10911091

adults from Country C were asked if human activity contributes to global warming. Complete parts​ (a), (b), and​ (c).
Adults who say that human activity contributes to global warming

Country Upper ACountry A

6666​%

Country Upper BCountry B

8989​%

Country Upper CCountry C

9595​%

​(a) Construct a

9999​%

confidence interval for the proportion of adults from

Country Upper ACountry A

who say human activity contributes to global warming.

left parenthesis nothing comma nothing right parenthesis0.6360.636,0.6840.684

​(Round to three decimal places as​ needed.)​(b) Construct a

9999​%

confidence interval for the proportion of adults from

Country Upper Country B

who say human activity contributes to global warming.

left parenthesis nothing comma nothing right parenthesis,

​(Round to three decimal places as​ needed.)
Answer #1

Ans:

A)

lower limit=0.66-2.576*SQRT(0.66*(1-0.66)/2600)=0.636

Upper limit=0.66+2.576*SQRT(0.66*(1-0.66)/2600)=0.684

B)

lower limit=0.89-2.576*SQRT(0.89*(1-0.89)/1108)=0.866

Upper limit=0.89+2.576*SQRT(0.89*(1-0.89)/1108)=0.914

C)

lower limit=0.95-2.576*SQRT(0.95*(1-0.95)/1091)=0.933

Upper limit=0.95+2.576*SQRT(0.95*(1-0.95)/1091)=0.967